OpenStudy (anonymous):

Parking in a student lot costs $2 for the first half hour and $1.75 for each hour thereafter. A partial hour is charged the same as a full hour. What is the longest time that a student can park in this lot for $9?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

its 1+(9-2)/1.75=1+4=5

OpenStudy (anonymous):

First 30 mins cost $2 X is the time after 30 mins We have equation: 2+1.75x=9 1.75x=7 x=4 -------> $9 can part: 4hours and 30 mins